OpenStudy (anonymous):

Prove by induction that if \[0 < a < 1\]then \[(1-a)^n \geq 1-na\]

OpenStudy (anonymous):

I think you know how to do it

OpenStudy (anonymous):

Let P(n) be the statement \[(1-a)^n \geq 1 - na; \ \ \ 0 < a < 1\]

OpenStudy (anonymous):

P(1) is the statement \[(1-a)^1 \geq 1 - (1)a\]iff\[1-a \geq 1-a\]which is true

OpenStudy (alfie):

Now prove that given a Pk true... Pk+1 is true as well! :)

OpenStudy (anonymous):

If all of P(1),..., P(k), are true, then in particular, P(k) is true. It must be shown that P(k+1) is true. P(k) is the statement \[(1-a)^k = 1-ka ; \ \ \ 0 < a < 1\]

OpenStudy (anonymous):

oops it's the staement \[(1-a)^k \geq 1 - ka ; \ \ \ 0 < a < 1\]

OpenStudy (anonymous):

let's multiply both sites by 1-a

OpenStudy (anonymous):

\[(1-a)^k \geq 1 - ka\]iff \[(1-a)^k(1-a)\geq(1-ka)(1-a)\]iff\[(1-a)^{k+1}\geq1 -a -ka+ka^2\]iff\[(1-a)^{k+1}\geq\]im stuck :(

OpenStudy (alfie):

ka^2 is an always positive term, how can that help you out?

OpenStudy (anonymous):

hmm.. it can be shown that (1-a)^{k+1} is bigger than the right hand side?

OpenStudy (alfie):

You agree that if you have... a > b+something positive => a > b ?

OpenStudy (anonymous):

yea \[(1-a)^{k+1}\geq1 -(k+1)a+ka^2\]

OpenStudy (anonymous):

so how do I end the proof

OpenStudy (anonymous):

since k is in N, and a^2 > 0, then ka^2 > 0, etc. etc.

OpenStudy (anonymous):

so I simply say \[(1-a)^{k+1}\geq1 -(k+1)a+ka^2\]iff\[(1-a)^{k+1}\geq1 -(k+1)a\]is that alright?

OpenStudy (alfie):

You wanted to prove that... (1-a)^k >= 1-ka. Now, since ka^2 is always positive, you can just take it... yeah, you got it ;)

OpenStudy (anonymous):

... which proves that P(k+1) is true.

OpenStudy (alfie):

right!